this is not homework.. please show work and answer... i have to understand this stuff. thanks so much!!

1. Find the zeros of y = (x + 7)(x – 3)(x – 2).
Then write the polynomial in standard form.
2. Factor x3 – 2x2 – 15x. Find the relative maximum and relative minimum.
When submitting an answer in the form of an exponent, use a caret (^) to indicate the power. For example, 53 would be written as 5^3.
3. Write a polynomial function in standard form with zeros at –5, –4, and 3.

4. Find any multiple zeros of ƒ(x) = x4 – 25x2 and state the multiplicity.

Answer by edjones(8007)

1.
y = (x + 7)(x – 3)(x – 2)
Zeroes: x=-7, x=3, x=2
x^3+2x^2-29x+42
.
2.
x3 – 2x2 – 15x
=x(x^2-2x-15)
=x(x-5)(x+3)
.
3.
–5, –4, and 3
(x+5)(x+4)(x-3)
=x^3+6x^2-7x-60
.
4.
x4 – 25x2
=(x^2+5)(x^2-5)
x^2=5
x=sqrt(5), x=-sqrt(5)
If you know imaginary numbers:
x^2=-5
x=sqrt(-5)=sqrt(5)*i
x=-sqrt(-5)=-sqrt(5)*i
.
Ed